OpenStudy (anonymous):

Stephanie is 3 yrs older than andrew. Twelve years from now the sum of their ages will be 30 yrs more than Stephanie's age at present. what linear function may be used to solve the problem? (any ideas)?

OpenStudy (fifciol):

Let S be the age of Stephanie and A age of Andrew S=A+3 S+12+A+12=S+30

OpenStudy (anonymous):

So firstly we know that the difference between their ages is 3. Let 'S' denote stephany's aeg and 'A' denote Andrew's age. Then we can set up an equation like this:\[\bf A+3=S\]Now 12 years later - means that by then both will be 12 years - the sum of their ages will be 30 years greater than stephanie's age atm, which is just S. Hence we get an equation like this:\[\bf (A+12)+(S+12)=(A+12)+(A+15)=30+(A+3)\]\[\bf \implies 2A+27=A+33 \implies A=6\] Now you can find S by plugging that in to our original equation. @bless